When selecting two cards from a standard deck of cards and the first card is replaced, what are the events called?
Answer #1

(Events are called independent)

Defination:→

An independent event is an event that has no connection to another event’s chances of happening (or not happening). In other words, the event has no effect on the probability of another event occurring.

When two events are independent, one event does not influence the probability of another event.

Example:→

if you choose a card from a deck of 52 cards, your probability of getting a king is 4 out of 52. Mathematically, you can write it like this:
P(king) = (number of kings in a deck of cards / total number of cards in a deck = 4/52 = 1/13

If you replace the king and choose again (assuming the cards are shuffled), the events are independent. Your probability remains the same. Choosing a card over and over again would be an independent event , because each time you choose a card (a “trial” in probability) it’s a separate, non-connected event.